# Alpha Divergence Losses for Biometric Verification

### v1

@@ Abstract
Performance in face and speaker verification is largely driven by margin-based softmax losses like CosFace and ArcFace. Recently introduced $\alpha$-divergence loss functions offer a compelling alternative, particularly for their ability to induce sparse solutions (when $\alpha>1$). However, integrating an angular margin—crucial for verification tasks—is not straightforward. We find this integration can be achieved in at least two distinct ways: via the reference measure (prior probabilities) or via the logits (unnormalized log-likelihoods). In this paper, we explore both pathways, deriving two novel margin-based $\alpha$-divergence losses: Q-Margin (margin in the reference measure) and A3M (margin in the logits). We identify and address a critical training instability in A3M—caused by the interplay of penalized logits and sparsity—with a simple yet effective prototype re-initialization strategy. Our methods achieve significant performance gains on the challenging IJB-B and IJB-C face verification benchmarks. We demonstrate similarly strong performance in speaker verification on VoxCeleb. Crucially, our models significantly outperform strong baselines at low false acceptance rates (FAR). This capability is critical for practical high-security applications, such as banking authentication, where minimizing false authentications is paramount.
@@ 1 Introduction
Margin-based softmax losses such as CosFace [cosface] and ArcFace [arcface] have become the de facto standard in modern face and speaker recognition. They achieve state-of-the-art accuracy by incorporating margin penalties to the logit of the ground truth identity, which explicitly maximizes inter-class separation. Such methods rely on the softargmax operator, which transforms logits into probabilities evaluated by the cross-entropy loss or equivalently, the Kullback–Leibler (KL) divergence with a uniform prior [alpha-loss]. To date, margin penalties have been applied exclusively through geometric modifications to logits (e.g., $\cos(\theta+m)$ in ArcFace, $\cos(\theta)-m$ in CosFace). While effective, these geometric margins are heuristic extensions of softmax that remain decoupled from its probabilistic formulation.
@@
Among the $f$-divergence family, the $\alpha$-divergence is particularly useful for biometric verification due to its tunable trade-off between smooth and sparse probability mappings. It defines a continuum of generalized entropy measures: recovering the Shannon negentropy (standard cross-entropy) as $\alpha\rightarrow 1$, the Gini negentropy (sparsemax [sparsemax]) at $\alpha=2$, and exhibiting intermediate behaviors for values such as $\alpha=1.5$. For $\alpha>1$, the $\alpha$-softargmax suppresses small probabilities, producing sparse posteriors that emphasize confident predictions, which in practice encourages more discriminative embeddings and sharper decision boundaries in face and speaker recognition.
However, naively combining $\alpha$-divergence with geometric margin modifications reveals a critical challenge: the loss’s sparsity-inducing gradients can force the ground-truth class probability to zero ($p_{y}=0$) during early training when prototypes are suboptimal, effectively eliminating the learning signal for the target class. This occurs because the $\cdot$$_{+}$ operator in the gradient formulation (defined later in the paper) zeroes out probabilities when logits fall below a threshold, interacting catastrophically with margin penalties that intentionally reduce target logits. Importantly, this incompatibility arises primarily during the early stages of training, when identity prototypes are poorly aligned and target logits are small. Once prototypes are well-formed, the gradients remain stable even under geometric margins—a property exploited by our second method.
We address this challenge through two complementary approaches that successfully unify $\alpha$-divergence with margin-based learning. Our first method, Q-Margin, introduces a probabilistic margin by encoding the penalty directly in the reference measure $\mathbf{q}$. By down-weighting the ground-truth class while keeping other classes uniform, it avoids the $p_{y}=0$ instability at the loss level.
Our second method, A3M-I, retains the conventional geometric margin formulation but stabilizes training through a mid-training prototype re-initialization. After several epochs, we recompute normalized identity prototypes from feature embeddings and replace the classifier weights, realigning the decision boundaries when sparse gradients would otherwise cause collapse. Although not probabilistic, A3M-I complements Q-Margin by addressing optimization-level instability.
@@
  • We identify and characterize a fundamental incompatibility between $\alpha$-divergence losses and geometric margin penalties, where sparsity-inducing gradients cause training collapse ($p_{y}=0$).
  • We propose Q-Margin, which introduces a probabilistic margin through non-uniform reference measures ($q_{y}=\exp(-s\cdot m)$), resolving the instability at the loss level and recovering CosFace as $\alpha\rightarrow 1$.
  • We propose A3M-I, which stabilizes training through mid-training prototype re-initialization, addressing optimization-level collapse for large and noisy datasets such as WebFace42M.
@@ 3.2 $\boldsymbol{f}$- and $\boldsymbol{\alpha}$-divergence Loss Framework
This framework generalizes the standard Cross-Entropy loss (recovered as $\alpha\to 1$ with $\textbf{q}=\textbf{1}$, yielding $f(u)=u\log u$) and allows for controllable sparsity for $\alpha>1$. We refer the reader to [alpha-loss] for the full derivation.
@@ 3.4.2 Gradients and Posterior Sparsity
where $\tau^{*}$ is a scalar value, computed iteratively, that ensures the probabilities sum to one $\left(\sum_{j}p_{j}=1\right)$. For $\alpha>1$, the $\cdot]_{+}=\max(0,\cdot)$ operator is active, meaning that if a logit $\theta_{j}$ falls below a certain threshold relative to $\tau^{*}$, the posterior probability $p_{j}$ becomes exactly 0. This mechanism is the source of the extreme sparsity observed in our experiments.
@@ 3.4.3 Efficient Computation of $\boldsymbol{\alpha}$-$\boldsymbol{\operatorname{softargmax}}$
Here, $f$ is the convex function generating the $\alpha$-divergence (Eq. 6), $f^{\prime}(u)=\frac{u^{\alpha-1}-1}{\alpha-1}$, $\alpha\neq 1$, the derivative of $f$, $f^{*}(u)=\frac{1}{\alpha}([1+(\alpha-1)u]^{\frac{\alpha}{\alpha-1}}_{+}-1)$, $\alpha\in(1,+\infty)$ is the convex conjugate of $f$, $f_{*}^{\prime}=(f^{*})^{\prime}$ the derivative of the convex conjugate (i.e., $\operatorname{softargmax}_{f}$). The components $q_{j}$ are from the reference measure and $\theta_{j}$ are the input logits.
@@ 3.4.5 A3M-I: Stabilizing A3M with Prototype Re-Initialization
To address the optimization collapse ($p_{y}=0$) detailed in Section 3.4.2, our method A3M-I enhances the A3M loss by incorporating a mid-training prototype weight initialization step. This procedure aims to replace potentially suboptimal prototypes, which, due to the sparsity-inducing nature of the loss, may have never received a gradient update with more optimal ones derived directly from the feature space.
@@ 4.3 Results
We first provide a quantitative analysis in Table 1 of the $p_{y}=0$ collapse and the resulting posterior sparsity. The A3M loss, which combines a geometric margin with $\alpha=1.25$, suffers from prototype collapse: 24.12% of all training images produced a zero probability for the ground-truth class, and 5.67% of all identities had every image of the identity produce a zero probability for the ground-truth class. Our proposed methods virtually eliminate this problem. A3M-I reduces the image collapse to 1.52% while Q-Margin’s probabilistic margin almost completely solves it (0.15%). Crucially, this stability is achieved while preserving the primary benefit of $\alpha$-divergence loss, as all methods produce an output vector that is $99.97\%$ sparse.
However, the final column reveals a fundamental behavioral difference. A3M-I still produces hard one-hot predictions for 27.44% of images. In contrast, Q-Margin does so for only 0.14% of images. This suggests that Q-Margin’s probabilistic margin acts as a softer regularizer that prevents over-confidence, whereas A3M-I retains the hard decision-making characteristics of the base A3M method.
Table 1: Sparsity and ground-truth collapse statistics for R-50 models on MS1MV3. We compare the $p_{y}=0$ collapse against the overall posterior sparsity, demonstrating that our methods address the collapse while retaining the desired sparsity of $\alpha$-divergence.
$p_{y}=0$ Collapse (%) Sparsity Statistics (%)
@@ 4.3.1 Ablation
To isolate the contribution of each component, we conducted two main ablations, detailed in Table 2. To analyze the impact of A3M-I’s mid-training prototype re-initialization, we compared its performance against the base A3M. The results show this technique is critically scale-dependent: on the smaller MS1MV3 dataset, re-initialization offered no clear advantage. However, this trend reversed on the massive WebFace42M dataset. Here, the A3M method still outperformed ArcFace and CosFace, but the full A3M-I method, with re-initialization, proved to be a critical component for top performance. This validates our hypothesis that this technique is essential for tackling the optimization challenges of large-scale noisy data.
To complete this study, we also applied the re-initialization technique to Q-Margin (termed Q-Margin-I in Table 2). This failed to improve performance on the IJB datasets, suggesting that the probabilistic margin of Q-Margin does not suffer from the same instability associated with the geometric margin of A3M.
@@ 5.3 Results
Both Q-Margin and A3M yield strong results at the two operating points of interest. However, to provide a broader assessment, we also report the average relative improvement over the full FAR interval between $10^{-3}$ and $10^{-4}$ FAR. We observe a significant improvement from Q-Margin ($\alpha=1.5,s=10,m=0.1$) over both baselines. Its performance is closely followed by that of A3M-I ($\alpha=1.75,s=5,m=0.1$).
@@ 6 Conclusions
Margin-based loss functions have driven the success of embedding-based networks in face recognition and speaker verification, while recent work on generalized cross-entropy ($\alpha$-divergence losses) has enabled principled control of sparsity in probability distributions and class-dependent penalization through the reference measure q. We propose a synergistic combination of these two paradigms and resolve a core challenge: sparsity-inducing gradients interact destructively with geometric margins, causing the ground-truth probability to collapse ($p_{y}=0$). Our method bridges information-theoretic loss design with geometric margin techniques.
Our first approach, Q-Margin, introduces a probabilistic margin by encoding margin penalties directly into a non-uniform q, avoiding gradient instability and offering strong robustness on large, noisy datasets. Our second method, A3M-I, provides an optimization-level remedy that preserves the standard geometric margin but prevents convergence collapse via mid-training prototype re-initialization using data-derived representations.
We targeted low-FAR operating regimes and show that both methods consistently outperform strong ArcFace and CosFace baselines on challenging face (IJB-B, IJB-C) and speaker (VoxCeleb1-H) benchmarks. Ablations further reveal practical hyperparameter settings and show that optimal $\alpha$ values are modality-dependent.
While A3M-I currently relies on geometric margins and prototype re-initialization, future work will explore theoretical formulations that avoid explicit geometric margins and integration with methods that adaptively adjust margins. In our framework, it would correspond to dynamically updating the non-uniform reference measure.

### v3

@@ Abstract
Performance in face and speaker verification is largely driven by margin-based softmax losses such as CosFace and ArcFace. Recently introduced $\alpha$-divergence loss functions offer a compelling alternative, particularly due to their ability to induce sparse solutions (when $\alpha>1$). However, integrating an angular margin—crucial for verification tasks—is not straightforward. We find that this integration can be achieved in at least two distinct ways: via the reference measure (prior probabilities) or via the logits (unnormalized log-likelihoods). In this paper, we explore both pathways, deriving two novel margin-based $\alpha$-divergence losses: Q-Margin (margin in the reference measure) and A3M (margin in the logits). We identify and address a training instability in A3M—caused by sparsity—with a simple yet effective prototype re-initialization strategy. Our methods achieve significant performance gains on the challenging IJB-B and IJB-C face verification benchmarks. We demonstrate similarly strong performance in speaker verification on VoxCeleb. Crucially, our models significantly outperform strong baselines at low false acceptance rates (FAR). This capability is critical for practical high-security applications, such as banking authentication, where minimizing false authentications is paramount. Finally, the sparsity of the $\alpha$-divergence-based posteriors enables memory-efficient training, which is crucial for datasets with millions of identities.
@@ 1 Introduction
Margin-based softmax losses such as CosFace [cosface] and ArcFace [arcface] have become the de facto standard in modern face and speaker recognition. They achieve state-of-the-art accuracy by incorporating margin penalties to the logit of the ground truth identity, which explicitly maximizes inter-class separation. Such methods rely on the $\operatorname{softargmax}$ operator, which transforms logits into probabilities evaluated by the cross-entropy loss or equivalently, the Kullback–Leibler (KL) divergence with a uniform prior [alpha-loss]. To date, margin penalties have been applied exclusively through geometric modifications to logits (e.g., $\cos(\theta+m)$ in ArcFace, $\cos(\theta)-m$ in CosFace). While effective, these geometric margins are heuristic extensions of softmax that remain decoupled from its probabilistic formulation.
@@
Among the $f$-divergence family, the $\alpha$-divergence is particularly useful for biometric verification due to its tunable trade-off between smooth and sparse probability mappings. It defines a continuum of generalized entropy measures: recovering the Shannon negentropy (standard cross-entropy) as $\alpha\rightarrow 1$, the Gini negentropy (sparsemax [sparsemax]) at $\alpha=2$, and exhibiting intermediate behaviors for values such as $\alpha=1.5$. For $\alpha>1$, the $\alpha$-$\operatorname{softargmax}$ suppresses small probabilities, producing sparse posteriors that emphasize confident predictions, which in practice encourages more discriminative embeddings and sharper decision boundaries in face and speaker recognition.
However, combining $\alpha$-divergence with margin-losses can be challenging: the loss’s sparsity-inducing gradients can effectively eliminate the learning signal for certain identities, such as those with few training examples. This occurs because the $\cdot$$_{+}$ operator in the gradient formulation (defined later in the paper) zeroes out probabilities when logits fall below a threshold. This can result in misalignment between certain prototypes and embeddings, resulting in $p_{y}=0$, where $y$ the ground-truth identity. Importantly, this misalignment arises primarily during the early stages of training, when identity prototypes are poorly aligned and target logits are small. Once prototypes are well-formed, the gradients remain stable even under geometric margins—a property exploited by our second method.
We address this challenge through two approaches that successfully unify $\alpha$-divergence with margin-based learning. Our first method, Q-Margin, introduces a probabilistic margin by encoding the penalty directly in the reference measure $\mathbf{q}$. By down-weighting the ground-truth class while keeping other classes uniform, it avoids the instability at the loss level.
Our second method, A3M-I, retains the conventional geometric margin formulation but stabilizes training through a mid-training prototype re-initialization. After several epochs, we recompute normalized identity prototypes from feature embeddings and replace the classifier weights, realigning the decision boundaries when sparse gradients would otherwise cause misalignment. Although not probabilistic, A3M-I complements Q-Margin by addressing optimization-level instability.
@@
  • We propose $\alpha$-divergence losses with margin penalties as an inherently-sparse alternative to standard logistic loss, for training highly discriminative identity embeddings.
  • We propose Q-Margin, which introduces a probabilistic margin through non-uniform reference measures ($q_{y}=\exp(-s\cdot m)$), which recovers CosFace as $\alpha\rightarrow 1$.
  • We propose A3M-I, which stabilizes training through mid-training prototype re-initialization, addressing optimization-level misalignment for large and noisy datasets such as WebFace42M.
@@ 3.2 $\boldsymbol{f}$- and $\boldsymbol{\alpha}$-divergence Loss Framework
This framework generalizes the standard Cross-Entropy loss (recovered as $\alpha\to 1$ with $\textbf{q}=\textbf{1}$, yielding $f(u)=u\log u-(u-1)$) and allows for controllable sparsity for $\alpha>1$. We refer the reader to [alpha-loss] for the full derivation.
@@ 3.4.2 Gradients and Posterior Sparsity
where $\tau^{*}$ is a scalar value, computed iteratively, that ensures the probabilities sum to one $\left(\sum_{j}p_{j}=1\right)$. For $\alpha>1$, the $\cdot]_{+}=\max(0,\cdot)$ operator is active, meaning that if a logit $\theta_{j}$ falls below a certain threshold relative to $\tau^{*}$, the posterior probability $p_{j}$ becomes exactly 0, resulting in sparse posterior distributions.
@@ 3.4.3 Efficient Computation of $\boldsymbol{\alpha}$-$\boldsymbol{\operatorname{softargmax}}$
Here, $f$ is the convex function generating the $\alpha$-divergence (Eq. 6), $f^{\prime}(u)=\frac{u^{\alpha-1}-1}{\alpha-1}$, $\alpha\neq 1$, the derivative of $f$, $f^{*}(v)=\frac{1}{\alpha}([1+(\alpha-1)v]^{\frac{\alpha}{\alpha-1}}_{+}-1)$, $\alpha\in(1,+\infty)$ is the convex conjugate of $f$ (i.e., $\operatorname{softmax}_{f}$), $f_{*}^{\prime}=(f^{*})^{\prime}$ the derivative of the convex conjugate (i.e., $\operatorname{softargmax}_{f}$). The components $q_{j}$ are from the reference measure and $\theta_{j}$ are the input logits.
@@ 3.4.5 A3M-I: Stabilizing A3M with Prototype Re-Initialization
To address the optimization instability detailed in Section 3.4.2, our method A3M-I enhances the A3M loss by incorporating a mid-training prototype weight initialization step. This procedure aims to replace potentially suboptimal prototypes, which, due to the sparsity-inducing nature of the loss, have been updated very rarely and are not well aligned with the embedding extractor, with more optimal ones derived directly from the feature space.
@@ 4.3 Results
We first provide a quantitative analysis in Table 1 of the misalignment between prototypes and embeddings (measured by identities and examples for which $p_{y}=0$) and the resulting posterior sparsity. The A3M loss, which combines a geometric margin $m=0.5$ with $\alpha=1.25$, suffers from severe misalignment: 24.12% of all training images produced a zero probability for the ground-truth class, and 5.67% of all identities had every image of the identity produce a zero probability for the ground-truth class. Our proposed methods virtually eliminate this problem. A3M-I (also with $m=0.5$) reduces the image-level misalignment to 1.52% while Q-Margin’s probabilistic margin almost completely solves it (0.15%).
Crucially, this stability is achieved while preserving the primary benefit of $\alpha$-divergence loss, as all methods produce an output vector that is $99.97\%$ sparse. Ideally, this extreme sparsity allows for the construction of a memory-efficient training framework where logit evaluation is constrained to neighboring prototypes. Such an approach, facilitated by sparse embedding layers and optimizers, would yield significant computational and memory gains without resorting to approximate methods.
The final column reveals another important behavioral difference between the methods. A3M-I still produces hard one-hot predictions for 27.44% of images. In contrast, Q-Margin does so for only 0.14% of images. This suggests that Q-Margin’s probabilistic margin acts as a softer regularizer that prevents over-confidence, whereas A3M-I retains the hard decision-making characteristics of the base A3M method.
Table 1: Statistics of sparsity and prototype misalignment of R-50 models trained on MS1MV3. We compare misalignment of prototypes (using $p_{y}=0$ as a cue) against the overall posterior sparsity, demonstrating that our methods address the misalignment while retaining the desired sparsity of $\alpha$-divergence. The statistics are collected after the final epoch.
Misalignment $p_{y}=0$ (%) Sparsity Statistics (%)
@@ 4.3.1 Ablation
To isolate the contribution of each component, we conducted two main ablations, detailed in Table 2. To analyze the impact of A3M-I’s mid-training prototype re-initialization, we compared its performance against the base A3M. The results show this technique is critically scale-dependent: on the smaller MS1MV3 dataset, re-initialization offered no clear advantage. However, this trend reversed on the massive WebFace42M dataset. Here, the A3M method still outperformed ArcFace and CosFace, but the full A3M-I method, with re-initialization, proved to be a critical component for top performance. This validates our hypothesis that this technique is helpful in tackling the optimization challenges of large-scale noisy data. To complete this study, we also applied the re-initialization technique to Q-Margin (termed Q-Margin-I in Table 2). This failed to improve performance on the IJB datasets, suggesting that the probabilistic margin of Q-Margin does not suffer from the same instability associated with the geometric margin of A3M.
@@ 5.3 Results
Both Q-Margin and A3M yield strong results at the two operating points of interest. To provide a broader assessment, we also report the average relative improvement over the interval between $10^{-3}$ and $10^{-4}$ FAR. We observe a significant improvement from Q-Margin ($\alpha=1.5,s=10,m=0.1$) over both baselines. Its performance is closely followed by that of A3M-I ($\alpha=1.75,s=5,m=0.1$).
@@ 6 Conclusions
Margin-based loss functions have driven the success of embedding-based networks in face recognition and speaker verification, while recent work on generalized cross-entropy ($\alpha$-divergence losses) has enabled principled control of sparsity in probability distributions and class-dependent penalization through the reference measure q. We propose a synergistic combination of these two paradigms and resolve a core challenge: sparsity-inducing gradients interact destructively with geometric margins, causing misaligment between prototypes and embeddings. Our method bridges information-theoretic loss design with geometric margin techniques.
Our first approach, Q-Margin, introduces a probabilistic margin by encoding margin penalties directly into a non-uniform q, avoiding gradient instability and offering strong robustness on large, noisy datasets. Our second method, A3M-I, provides an optimization-level remedy that preserves the standard geometric margin but mitigates misalignment via mid-training prototype re-initialization using data-derived representations.
We targeted low-FAR operating regimes and show that both methods consistently outperform strong ArcFace and CosFace baselines on challenging face (IJB-B, IJB-C) and speaker (VoxCeleb1-H) benchmarks. Ablations also reveal practical hyperparameter settings and show that optimal $\alpha$ values are modality-dependent.
While A3M-I currently relies on geometric margins and prototype re-initialization, future work will explore theoretical formulations that avoid explicit geometric margins and integration with methods that adaptively adjust margins. In our framework, it would correspond to dynamically updating the non-uniform reference measure. Finally, we plan to exploit the inherent sparsity of the proposed losses towards a memory-efficient implementation.